OpenStudy (anonymous):

MEDAL AND FAN FOR ALGEBRA 1 QUESTIONS. Harris has two pipes to fill his aquarium. Pipe A can fill the aquarium at the rate of 4 gallons per minute. Pipe B can fill the aquarium at the rate of 840 liters per hour. [1 gallon = 3.79 liters]

OpenStudy (anonymous):

Part A: Which pipe will fill the aquarium faster? Justify your answer by comparing the rates after converting to the same units. (5 points) Part B: If the total volume of the aquarium was 500 gallons and you checked on how full it was every 0.25 hours, describe appropriate scales on the x-axis and y-axis to graph the rate at which the faster pipe fills the aquarium. (5 points)

OpenStudy (anonymous):

Okay for part 1, the answer will be pipe A. 4*3.79 = 15.16 liters/minute 15.16*60 = 909.6 liters/hour

OpenStudy (anonymous):

omg thank you so much. Can you help me with Part B too ?

OpenStudy (anonymous):

Okay, I'll see if I can :)

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Okay, so we're given: total volume of tank = 500 gallons check every 0.25 hours = 15 minutes Pipe A = 909.6 liters/hour

OpenStudy (anonymous):

ok ill help u hold up a second

OpenStudy (anonymous):

500*3.79 = 1895 liters 909.6/4 = 227.4 liters/minute

OpenStudy (anonymous):

1895/227.4 = 8.333... So this means it takes a little over 8 minutes to fill the tank

OpenStudy (anonymous):

Is it asking for the answer or is it asking for what the graph would look like ?

OpenStudy (anonymous):

What the graph would look like, since it's asking for the scale

OpenStudy (anonymous):

So your x-axis should go by a scale of 1 minute and your y-axis should go by a scale of about 227 liters

OpenStudy (anonymous):

or would it be 20 30 60 for the y axis and 15 minutes for the x axis

OpenStudy (anonymous):

Where did the 20 30 60 come from?

OpenStudy (anonymous):

I dont even know im so confused lol

OpenStudy (anonymous):

If you plot on a scale of 15 minutes, you'd only have one point

OpenStudy (anonymous):

15 , 30 , 45

OpenStudy (anonymous):

hope this helps

OpenStudy (anonymous):

okay let's just try the next question lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Winnie poured 14 cups of water into a rectangular container measuring 13 inches by 7 inches by 6 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. (3 points) Part B: How many cubic inches of water was poured into the container? (3 points) Part C: What was the height of the water in the container? Show your work. (4 points)

OpenStudy (anonymous):



OpenStudy (anonymous):

so The answer is 7.6inches

OpenStudy (anonymous):

can i get a medal?

OpenStudy (anonymous):

Zoe why don't you try doing your own work instead of copying...your teachers are going to catch you